Brief Exercise 1-8 Use the basic accounting equation to answer these questions. (a) The liabilities of Jantz Company are $95,860 and the stockholders equity is $269,500. What is the amount of Jantz Companys total assets? Total assets (b) The total assets of Foley Company are $189,900 and its stockholders equity is $85,300. What is the amount of its total liabilities? otel labilities (c) The total assets of Sundberg Co. are $947,700 and its liabilities are equal to one-fourth of its total assets. What is the amount of Sundberg Co.s stockholders equity? Stockholders equity Click if you would like to Show Work for this question: Open Show Work Question Attempts:0 of 3 used SAVE FOR LATER SUBMET ANSWER
Answer #1

a) Total assets = Total liabilities + Total Stockholder's equity

= 95860+269500

Total assets = 365360

b) Total liabilities = Total assets - Total stockholder's equity

= 189900-85300

Total liabilities = 104600

c) Total stockholders equity = Total assets -Total liabilities

=947700- (947700/4)

Total Stockholder's equity = 710775